# Quantifying and Mitigating Privacy Risks for Tabular Generative Models

### v1

@@
Abstract.
Synthetic data from generative models emerges as the privacy-preserving data sharing solution. Such a synthetic data set shall resemble the original data without revealing identifiable private information. The backbone technology of tabular synthesizers is rooted in image generative models, ranging from Generative Adversarial Networks (GANs) to recent diffusion models. Recent prior work sheds light on the utility-privacy tradeoff on tabular data, revealing and quantifying privacy risks on synthetic data. However, the focus is limited to a small number of privacy attacks and tabular synthesizers, particularly GAN-based, and overlooks membership inference attacks and defense strategies, i.e., differential privacy. To bridge the gap, we address two research questions: (i) which type of tabular generative models can achieve better utility-privacy tradeoff, considering both a broader collection of synthesizers and their performance against membership inference attacks; (ii) what additional privacy guarantee can be achieved by means of the differentially private stochastic gradient descent algorithm (DP-SGD). We first conduct an exhaustive empirical analysis, highlighting the utility-privacy tradeoff of five state-of-the-art tabular synthesizers, against eight privacy attacks, with a special focus on membership inference attacks. Motivated by the observation of high data quality but also high privacy risk in tabular diffusion, we propose DP-TLDM, Differentially Private Tabular Latent Diffusion Model, which is composed of an autoencoder network to encode the tabular data and a latent diffusion model to synthesize the latent tables. Following the emerging $f$-DP framework, we apply DP-SGD to train the auto-encoder in combination with batch clipping and use the separation value as the privacy metric to better capture the privacy gain from DP algorithms. Our empirical evaluation demonstrates that DP-TLDM is capable of achieving a meaningful theoretical privacy guarantee while also significantly enhancing the utility of synthetic data. Specifically, compared to other DP-protected tabular generative models, DP-TLDM improves the synthetic quality by an average of 35% in data resemblance, 15% in the utility for downstream tasks, and 50% in data discriminability, all while preserving a comparable level of privacy risk.
@@ Abstract.
journalyear: YYYYjournalvolume: YYYYjournalnumber: Xdoi: XXXXXXX.XXXXXXX
1. Introduction
@@
High-quality synthetic data obtained from generative models are increasingly used to augment and substitute real data, boosting data utility for individuals and enterprises (Ali et al., 2022; Chambon et al., 2022a, b). As synthetic data resembles real data, it can be used to accelerate data-driven knowledge discovery and still abide by data protection regulations, e.g., GDPR (eu2, 2016), which restricts the collection and accessibility of real data. A key requirement for the adoption of these models in the industry is their ability to preserve the privacy of the real data (Yoon et al., 2020; Giomi et al., 2023). Consider for instance medical institutes that own a subset of patients’ data that cannot be shared freely and are subject to lengthy regulatory auditing. Alternatively, through a trusted party that first trains the generative model, a complete set of patients’ synthetic data can be generated and distributed to all institutes that in turn design their own medical analysis based on these data.
While the focus of generative models lies on producing synthetic data highly similar to and indiscernible from the real data, a rising concern is the real data privacy leakage caused by the synthetic data (Carlini et al., 2023; Hu and Pang, 2023; Duan et al., 2023). These studies highlight privacy vulnerabilities associated with synthetic data across specific domains, especially in image processing, and pertain to various generative models, including Bayesian networks, generative adversarial networks (GANs), and, more recently, diffusion processes. These privacy risks are materialized in attacks that are able to obtain training data, under various assumptions on the availability of model knowledge, i.e., white-box v.s. black-box attacks.
@@
In the tabular data domain, Anonymeter (Giomi et al., 2023) is the first framework that focuses on the privacy and utility trade-off of synthetic tables and introduces three privacy attacks relevant to tabular data: i) singling out attacks that identify individuals through unique combinations of attributes in the synthetic dataset; ii) linkability attacks that associate two or more records by searching through neighbors in the synthetic dataset; and iii) attribute inference attacks that deduce undisclosed attribute values through the synthetic dataset. While it sheds light on quantifying the privacy-utility trade-off, (Giomi et al., 2023) focuses on tabular GAN models, leaving the question of how this tradeoff behaves for different tabular generative models unaddressed. More importantly, the critically important category of Membership Inference Attacks (MIA) (Hayes et al., 2019; Chen et al., 2020b; Mukherjee et al., 2021), which present stronger adversarial assumptions and infer whether specific data records are present in the training set, is overlooked. Last but not least, the impact of adopting privacy-enhancing strategies, such as differential privacy, on synthetic tabular data is largely unexplored by prior art.
@@
Differential privacy (DP) (Dwork et al., 2006) has received much attention as a solution to the problem of preserving individual privacy when releasing data. To incorporate DP in the training of deep neural models with stochastic gradient descent (SGD), DP-SGD (Abadi et al., 2016) obfuscates gradient updates by adding calibrated statistical noise that is controlled by a privacy budget. There are two main DP analysis frameworks, ($\epsilon,\delta$)-DP (Dwork et al., 2006), and emerging f-DP (Dong et al., 2021), where the former uses $\epsilon$ to define the privacy budget and the latter uses the separation value, which is the distance between the actual trade-off function of false positive and false negative and the ideal one, where no privacy leaks. A smaller privacy budget or separation value leads to adding more obfuscation noise to the gradients, degrading the performance of the underlying models. It is a long-standing challenge to apply meaningful $\epsilon$ or separation value while achieving satisfactory learning outcomes for image classification (Abadi et al., 2016) and synthesizing (Jordon et al., 2018). The privacy enhancement of DP on tabular generative models is yet to be explored, especially concerning different genres of generative models.
We recognize two research gaps between the need to use synthetic data and current solutions: (i) which type of tabular generative models provide an effective privacy-utility trade-off, e.g., GANs v.s. diffusion models, (ii) how to apply and parameterize DP-SGD to enhance the privacy guarantee of synthetic tables. In this paper, we address them by conducting exhaustive empirical analyses to quantify and improve the privacy-utility trade-off for tabular generative models. We consider five tabular synthesizers, covering Gaussian Copulas, CopulaGAN, CTGAN, ADS-GAN, and tabular diffusion models (TabDDPM). Our evaluation metrics on the utility of synthetic data include data resemblance, discriminability, and downstream utility. As for the privacy risk, we consider four types of attacks, namely singling out, linkability, attribute inference (AIA), and membership inference attacks (MIA). Among those attacks, MIA employs stronger adversarial assumptions, leveraging the knowledge of generative models and their training data sets. To gain an in-depth analysis of MIA under different adversarial knowledge, we consider 5 types of MIA attacks, amounting to eight attacks evaluted in this study. Our extensive evaluation on four data sets highlights that synthetic data from the tabular diffusion model achieves remarkable data quality, compared to three other GAN-based synthesizers. In terms of privacy risk, TabDDPM can achieve satisfying performance against singling out, linkability, but suffers against attribute and especially membership inference attacks.
@@
Based on the insights from our empirical study, we propose a novel differentially private latent tabular diffusion model, DP-TLDM, composed of an autoencoder and a diffusion model. Different from the existing TabDDPM, we first encode the tabular data into a continuous latent space, using the autoencoder network. This brings the advantage of a unified and compact representation of categorical variables, in contrast to the typical one-hot encoding. We then use the latent representation as input to the backbone diffusion model, which captures the data synthesis as a sequence of denoising processes (Kotelnikov et al., 2023). To guard the proposed latent tabular diffusion against privacy attacks, we train the auto-encoder using DP-SGD. We follow the $f$-DP framework (Dong et al., 2021), which provides a precise parameterization of DP-SGD, specifically through the separation measure — a metric quantifying the maximum difference between false positive and false negatives when comparing a random guess and DP-protected algorithm. Thanks to the post-processing guarantees of DP, the backbone latent diffusion training is also protected by the DP. We extensively evaluate the proposed DP-TLDM against DP-CTGAN and DP-TabDDPM, where DP-SGD is used to train CTGAN and TabDDPM, showing a remarkable performance — reducing the privacy risk especially against MIA while maintaining a significant high data utility compared to the other two synthesizers. We make the following concrete contributions:
  • Extensive and thorough empirical evaluation on the utility-privacy trade-off of five tabular synthesizers under eight privacy attacks, including the strong adversaries of no-box membership inference attacks.
  • Our key insight is that tabular diffusion is shown to achieve higher synthetic data quality, compared to the Gaussian copula model and GAN-based models. Also, it is able to withstand the singling out, linkability, and attribute inference attacks, but displays high privacy risks against MIA.
@@
  • We design DP-TLDM, a novel latent tabular diffusion model trained by DP-SGD that uses batch clipping on gradients and Gaussian noising mechanism. We follow the $f$-DP framework and adopt the theoretical separation value as the privacy metric.
@@ 2. Related Studies
In this section, we provide a general overview of the generative models, which are first designed and developed for image data. The specific tabular generative models are discussed in Section 3.
Generative models have gained extensive attention in machine learning, ranging from simple to deep generative architectures. Fundamental machine learning models, including the Gaussian Mixture Model (GMM) (McLachlan and Basford, 1988) and Naive Bayes (Gelman et al., 1995) utilize joint probability distributions to represent and sample the data. While these models have been amongst the earliest models used for data generation, they usually have limited expressiveness. Deep generative models, on the other hand, excel in capturing complex data patterns and are more widely used nowadays.
Generative Adversarial Networks (GANs) (Goodfellow et al., 2014; Arjovsky et al., 2017; Brock et al., 2018), typically consisting of a generator and a discriminator, are one of the most popular deep generative models. Flow-based generative models, also known as normalizing flows (Dinh et al., 2014; Xu et al., 2022) are another type of explicit model that employ a series of invertible transformation functions to convert a simple distribution into the inherently complex data distribution. Diffusion models (Sohl-Dickstein et al., 2015) are recently emergent generative models known for their impressive performance in various fields, including image synthesis (Ho et al., 2020; Karras et al., 2022), text-to-image generation (Rombach et al., 2022; Saharia et al., 2022; Ramesh et al., 2022), spatio-temporal data modeling (Tashiro et al., 2021) and so on. These models (Ho et al., 2020; Rombach et al., 2022; Karras et al., 2022) operate via a forward diffusion process, noising the original data, followed by a subsequent reverse process denoising the data.
Despite the impressive performance and application of deep generative models, recent works have also raised significant concerns regarding the potential privacy risks of these models. A vast body of related studies on privacy attacks can be categorized by various attack types, including (i) membership inference attacks (MIA) (Shokri et al., 2017; Carlini et al., 2023; Chen et al., 2020b; Hayes et al., 2019), inferring whether a certain data record is in the training set; (ii) attribute inference attacks (Stadler et al., 2020; Giomi et al., 2023), deducing sensitive attributes of the training data; (iii) replication attacks (Carlini et al., 2023; Hu and Pang, 2021; Somepalli et al., 2023), reproducing the training data or hidden generative models; (iv) adversarial attacks (Szegedy et al., 2014; Millière, 2022; Gao et al., 2023), deceiving generative models through crafted input data at the inference stage; and (v) backdoor attacks (Yang et al., 2017; Ding et al., 2019), inserting hidden vulnerabilities into models at the training stage.
Membership inference attacks can be further categorized into white-box, no-box and black-box attacks based on the availability of model information. In white-box attacks, where attackers have access to the internals of generators, several works (Carlini et al., 2023; Hu and Pang, 2023; Zhu et al., 2023) have proposed loss-based techniques for conducting MIA on diffusion models. Meanwhile, the gradient information from the generator can be utilized to perform MIA against GANs (Chen et al., 2020b; Liu et al., 2019). In the gray-box setting, adversaries possess partial control over the victim model. In this context, the latent code from the generator is adjusted by (Chen et al., 2020b) and (Liu et al., 2019) to attack GANs, while intermediate generative results obtained from different diffusion steps are leveraged to infer the membership of a query in diffusion models (Duan et al., 2023; Zhu et al., 2023). The most challenging and realistic scenario happens in the black-box setting where the prior knowledge of attackers is limited only to generated samples. One work (Hayes et al., 2019) trains shadow discriminators from GANs to output the confidence score of being a training sample. Other less computationally intensive attacks rely on the fluctuations of probabilities (Fu et al., 2023) or semantic distances (Zhu et al., 2023; Wu et al., 2022) between the query record and its neighbors to infer membership.
Privacy enhancing methodologies have been studied to address potential privacy risks. Simple defense strategies, such as fine-pruning (Alfeld et al., 2016; Ding et al., 2019), deduplicating training data (Carlini et al., 2023), and data augmentation (Duan et al., 2023) have been proposed to counteract malicious attacks. Despite their ease of implementation, these techniques do not exhibit a high level of effectiveness. More sophisticated strategies have been further developed for different generative models. For example, privGAN (Mukherjee et al., 2021), RoCGAN (Chrysos et al., 2020) and PATE-GAN (Jordon et al., 2018) equip GANs with privacy protection through strategic changes to the model architecture. While these strategies exhibit enhanced effectiveness, they remain model-specific in nature, thereby limiting their broader applicability across various generative model paradigms.
@@
Conversely, differential privacy (DP) has proven to be a more general and effective defense mechanism for preventing privacy leaks. DP-SGD and its variants have been widely adopted for privately training deep generative models. DPGAN (Xie et al., 2018) applies the DP-SGD algorithm directly to the discriminator component within GANs. In contrast, GS-WGAN (Chen et al., 2020a) implements DP-SGD on the gradients transferred from the discriminator to the generator. The utility of DP-SGD is also extended beyond GANs and applied in normalizing flows for tabular data synthesis (Waites and Cummings, 2021; Lee et al., 2022). Moreover, in the context of emerging diffusion models, adaptions of DP-SGD are considered as well. One study (Dockhorn et al., 2022) applied the classic DP-SGD algorithm with one modification involving sampling multiple time steps of a single data point when computing the loss. Building on this, another study (Ghalebikesabi et al., 2023) further presented the effectiveness of three other techniques, namely pre-training, augmentation multiplicity, and modified time step sampling. While DP-SGD is deemed a strong countermeasure for privacy leaks, it comes at the cost of sample quality and longer training times.
3. Risk-Utility Quantification
Figure 1. The risk-utility quantification and enhancing framework: from training to generation. Starting with the original tabular data, synthetic data is generated by synthesizers with or without DP protection. This synthetic data is then evaluated from two critical perspectives: utility (resemblance, discriminability, and utility) and privacy risk (singling out, linkability, attribute inference attack (AIA), and membership inference attack (MIA)).
3.1. Generative models
GAN-based models
We consider three different Generative Adversarial Networks (GAN) models. First, CTGAN (Xu et al., 2019) employs GANs with a focus on conditional generation. It employs mode-specific normalization to handle non-Gaussian distributions in continuous columns, and a conditional generator to address class imbalance in categorical columns. Second, CopulaGAN (Synthetic Data Vault, 2023) improves upon CTGAN by utilizing cumulative distribution function-based transformations with Gaussian Copulas. It also performs inference using a likelihood approach, enhancing CTGAN’s ability to learn real data trends. Third, ADS-GAN (Yoon et al., 2020) is a conditional GAN framework that generates synthetic data while minimizing re-identification risk. It achieves a certain degree of anonymization by incorporating a record-level identifiability metric into the generator’s loss function.
Statistical models
Here we consider Gaussian Copula models. In the Gaussian Copula (GC) method (Nelsen, 2006), the training data is used to obtain a Gaussian joint probability distribution that captures both marginal distributions and interdependence structures.
Difussion models
Diffusion models have recently become the leading paradigm in generative models for computer vision and NLP. In our framework, we consider TabDDPM (Kotelnikov et al., 2023), which extends diffusion models to tabular datasets, outperforming existing GAN/VAE alternatives. It employs the Gaussian diffusion process, a key component of the original DDPM (Ho et al., 2020), to model numerical columns effectively. It also uses the multinomial diffusion process to model categorical and binary features and to introduce uniform noise across classes to corrupt data.
3.2. Utility metrics
Resemblance
Discriminability
Utility
3.3. Threat model
In particular, for linkability attacks, the needed auxiliary data comprises two disjoint sets of attributes $T[:,A]$ and $T[:,B]$, extracted from $T$. For attribute inference attacks, the attacker needs to know the values of a set of attributes $T[:,A]$. In shadow modeling-based membership inference attacks, the attacker, alongside the target set $T$, has access to another reference dataset $X_{R}$. As in (Stadler et al., 2022), this dataset mirrors the distribution of the training dataset and may or may not have overlapping records with it.
3.4. Attacks
As mentioned before, to quantify the privacy risks of tabular data synthesizers, we employ four distinct attacks in our evaluation framework. Three of these attacks—singling out, linkability, and attribute inference attack (AIA)—are derived from the guidelines set by the European General Data Protection Regulation (GDPR), following (Giomi et al., 2023). Additionally, we incorporate the membership inference attack (MIA), a well-established but previously omitted facet in (Giomi et al., 2023). This enhances our privacy quantification framework by allowing for the measurement of an additional and widely acknowledged dimension of privacy in synthetic data.
The implementation of singling out, linkability and AIA follows the open-source library Anonymeter  [^3] https://github.com/statice/anonymeter from (Giomi et al., 2023) while the membership inference attack (MIA) follows the TAPAS toolbox  [^4] https://github.com/alan-turing-institute/tapas from (Houssiau et al., 2022).
Singling out
The singling out attack (Giomi et al., 2023) aims to create predicates from the synthetic dataset that could identify an individual present in the training dataset. For example, if an attacker can determine that a dataset has only one individual with attributes like age: 25, height: 168, weight: 62, and cholesterol: 1, that individual is considered ”singled out”.
Following the implementation in (Giomi et al., 2023), two algorithms are applied, the univariate algorithm and the multivariate algorithm. Both algorithms are based on the intuition that unique values or combinations of unique values in the synthetic data may also be unique in the original data. In the univariate algorithm, unique values are sampled for each attribute to obtain a random selection of predicates. In the multivariate algorithm, this is done for full records to obtain multivariate predicates.
Linkability
The linkability attack (Giomi et al., 2023) aims to associate two or more records. The linkability arises when an attacker has two disjoint sets of attributes $X_{A}$ (e.g., age and height) and $X_{B}$ (e.g., weight and cholesterol level) from the original dataset, so that it can use the synthetic dataset to determine that two records $x_{a}\in X_{A}$ and $x_{b}\in X_{B}$ belong to the same individual.
Attribute inference attack (AIA)
The AIA attack (Giomi et al., 2023) involves deducing undisclosed attribute values from information in the synthetic dataset. If an attacker knows certain attributes of an individual, such as age: 25, height: 168, and weight: 62, they might use the synthetic dataset to infer the cholesterol level of the individual.
Membership inference attack (MIA)
The MIA attack (Shokri et al., 2017) aims to determine if a specific data record is present in the training dataset. MIA attacks have gained substantial attention within the research community, leading to various proposed strategies for inferring the membership status of synthetic data points. In our framework, we employ three types of MIA strategies, based on shadow modeling, distance and probability. This diverse set accommodates a range of adversarial scenarios, recognizing different adversary capabilities and constraints.
In the shadow modeling approach, given a reference dataset $X_{R}$, which shares the same distribution as the training dataset, and a target record $x_{t}$, multiple training sets $X_{i}$ from $X_{R}$ are sampled and shadow models trained to generate synthetic datasets $X_{s}$ from $X_{i}$ and $X_{s}^{\prime}$ from $X_{i}^{\prime}=X_{i}\cup x_{t}$. A classifier is then trained on labeled synthetic datasets $X_{s}$ and $X_{s}^{\prime}$ to predict the presence of target records in the training data $X_{train}$. To reduce the effect of high-dimensionality and sampling uncertainty, instead of directly training on $X_{s}$ and $X_{s}^{\prime}$, the NaiveGroundhog strategy uses a basic feature set $F_{naive}$ for training, while HistGroundhog utilizes a histogram feature set $F_{hist}$ with marginal frequency counts of each data attribute (Stadler et al., 2022).
Metrics
Relative Risk Indicator The attacker success rate is a common metric for Membership Inference Attacks (MIA). However, for singling out, likability, and attribute inference attacks, as noted by (Giomi et al., 2023), there’s a distinction: certain information might be inferred from patterns inherent in the entire population $X_{ori}$ rather than solely from the training dataset $X_{train}$ and its synthetic counterpart.
Following (Giomi et al., 2023), the original dataset $X_{ori}$ is split into two disjoint partitions: $X_{train}$ and $X_{control}$. Privacy risk is then quantified by comparing attacker success rates for targets drawn from $X_{train}$ and $X_{control}$:
(1) $R=\frac{\hat{\tau}_{train}-\hat{\tau}_{control}}{1-\hat{\tau}_{control}}$
4. Empirical Analysis

### v2

@@
Abstract
Synthetic data from generative models emerges as the privacy-preserving data sharing solution. Such a synthetic data set shall resemble the original data without revealing identifiable private information. Till date, the prior focus on limited types of tabular synthesizers and small number of privacy attacks, particularly on Generative Adversarial Networks, and overlooks membership inference attacks and defense strategies, i.e., differential privacy. Motivated by the conundrum of keeping high data quality and low privacy risk of synthetic data tables, we propose DP-TLDM, Differentially Private Tabular Latent Diffusion Model, which is composed of an autoencoder network to encode the tabular data and a latent diffusion model to synthesize the latent tables. Following the emerging $f$-DP framework, we apply DP-SGD to train the auto-encoder in combination with batch clipping and use the separation value as the privacy metric to better capture the privacy gain from DP algorithms. Our empirical evaluation demonstrates that DP-TLDM is capable of achieving a meaningful theoretical privacy guarantee while also significantly enhancing the utility of synthetic data. Specifically, compared to other DP-protected tabular generative models, DP-TLDM improves the synthetic quality by an average of 35% in data resemblance, 15% in the utility for downstream tasks, and 50% in data discriminability, all while preserving a comparable level of privacy risk.
Keywords:
@@ Abstract
1 Introduction
@@
High-quality synthetic data obtained from generative models are increasingly used to augment and substitute real data, boosting data utility for individuals and enterprises [2, 6, 7]. As synthetic data resembles real data, it can be used to accelerate data-driven knowledge discovery and still abide by data protection regulations, e.g., GDPR [16], which restricts the collection and accessibility of real data. A key requirement for the adoption of these models in the industry is their ability to preserve the privacy of the real data [49, 19]. Consider for instance medical institutes that own a subset of patients’ data that cannot be shared freely and are subject to lengthy regulatory auditing. Alternatively, through a trusted party that first trains the generative model, a complete set of patients’ synthetic data can be generated and distributed to all institutes that in turn design their own medical analysis based on these data. While the focus of generative models lies on producing synthetic data highly similar to and indiscernible from the real data, a rising concern is the real data privacy leakage caused by the synthetic data [5, 27, 14]. These studies highlight privacy vulnerabilities associated with synthetic data across specific domains, especially in image processing, and pertain to various generative models, including Bayesian networks, generative adversarial networks (GANs), and, more recently, diffusion processes. These privacy risks are materialized in attacks that are able to obtain training data, under various assumptions on the availability of model knowledge, i.e., white-box v.s. black-box attacks.
@@
In the tabular data domain, Anonymeter [19] is the first framework that focuses on the privacy and utility trade-off of synthetic tables and introduces three privacy attacks relevant to tabular data: singling out attacks, linkability attacks, and attribute inference attacks. While it sheds light on quantifying the privacy-utility trade-off, [19] focuses on tabular GAN models, i.e., CTGANs [48], leaving the question of how this tradeoff behaves for different tabular generative models unaddressed. More importantly, the critically important category of Membership Inference Attacks (MIA) [21, 9, 33], which present stronger adversarial assumptions and infer whether specific data records are present in the training set, is overlooked. Last but not least, the impact of adopting privacy-enhancing strategies, such as differential privacy, on synthetic tabular data is largely unexplored by prior art.
@@
Differential privacy (DP) [15] has received much attention as a solution to the problem of preserving individual privacy when releasing data. To incorporate DP in the training of deep neural models with stochastic gradient descent (SGD), DP-SGD [1] obfuscates gradient updates by adding calibrated statistical noise that is controlled by a privacy budget. There are two main DP analysis frameworks, ($\epsilon,\delta$)-DP [15], and emerging f-DP [13], where the former uses $\epsilon$ to define the privacy budget and the latter uses the separation value, which is the distance between the actual trade-off function of false positive and false negative and the ideal one, where no privacy leaks. A smaller privacy budget or separation value leads to adding more obfuscation noise to the gradients, degrading the performance of the underlying models. It is a long-standing challenge to apply meaningful $\epsilon$ or separation value while achieving satisfactory learning outcomes for image classification [1] and synthesizing [28]. The privacy enhancement of DP on tabular generative models is yet to be explored, especially concerning different genres of generative models.
@@
Based on the insights from our empirical study, we propose a novel differentially private latent tabular diffusion model, DP-TLDM, composed of an autoencoder and a diffusion model. Different from the existing TabDDPM, we first encode the tabular data into a continuous latent space, using the autoencoder network. This brings the advantage of a unified and compact representation of categorical variables, in contrast to the typical one-hot encoding. We then use the latent representation as input to the backbone diffusion model, which captures the data synthesis as a sequence of denoising processes [30]. To guard the proposed latent tabular diffusion against privacy attacks, we train the auto-encoder using DP-SGD. We follow the $f$-DP framework [13], which provides a precise parameterization of DP-SGD, specifically through the separation measure — a metric quantifying the maximum difference between false positive and false negatives when comparing a random guess and DP-protected algorithm. Thanks to the post-processing guarantees of DP, the backbone latent diffusion training is also protected by the DP. We extensively evaluate the proposed DP-TLDM against DP-CTGAN and DP-TabDDPM, where DP-SGD is used to train CTGAN and TabDDPM, showing a remarkable performance — reducing the privacy risk especially against MIA while maintaining a significant high data utility compared to the other two synthesizers. We make the following concrete contributions:
@@
  • We design DP-TLDM, a novel latent tabular diffusion model trained by DP-SGD that uses batch clipping on gradients and Gaussian noising mechanism. Our approach leverages the $f$-DP framework, where we propose a new theoretical privacy metric, termed separation, to enhance privacy guarantees.
@@ 2 Related Studies
In this section, we provide a general overview of the generative models and privacy on the tabular data.
Tabular generative models Current state-of-the-art introduces several deep generative models for tabular data synthesis. TableGAN [36] implements an auxiliary classification model along with discriminator training to enhance column dependency in the synthetic data. CTGAN [48] improves data synthesis by introducing several preprocessing steps for categorical, continuous or mixed data types which encode data columns into suitable form for GAN training. The conditional vector designed by CTGAN [48] and later improved by Ctab-GAN [51] also helps the GAN training to reduce mode-collapse on minority categories. Drawback of these methods is also clear that there is loss of information during the transformation from table to latent vector. Therefore, GAN cannot learn the knowledge from the information that loses during this compression. TabDDPM [30] is based on denoising diffusion probabilistic models (DDPM) [22], it uses two different diffusion models to synthesize categorical and continuous columns. Privacy attacks Despite the impressive performance and application of deep generative models, recent works have also raised significant concerns regarding the potential privacy risks of these models. A vast body of related studies on privacy attacks can be categorized by various attack types, including (i) membership inference attacks (MIA) [38, 5, 9, 21], inferring whether a certain data record is in the training set; (ii) attribute inference attacks [20, 19], deducing sensitive attributes of the training data; (iii) replication attacks [5, 26, 39], reproducing the training data or hidden generative models; (iv) adversarial attacks [42, 32, 17], deceiving generative models through crafted input data at the inference stage. Membership inference attacks can be further categorized into white-box, no-box and black-box attacks based on the availability of model information. In white-box attacks, where attackers have access to the internals of generators, several works [5, 27, 52] have proposed loss-based techniques for conducting MIA on diffusion models. Black-box setting assumes the prior knowledge of attackers is limited only to generated samples [52, 46].
@@
Privacy enhancing methodologies have been studied to address potential privacy risks. DP-SGD and its variants have been widely adopted for privately training deep generative models. DPGAN [47] applies the DP-SGD algorithm directly to the discriminator component within GANs. In contrast, GS-WGAN [8] implements DP-SGD on the gradients transferred from the discriminator to the generator. The utility of DP-SGD is also extended beyond GANs and applied in normalizing flows for tabular data synthesis [44, 31]. Moreover, in the context of emerging diffusion models, adaptions of DP-SGD are considered as well. One study [12] applied the classic DP-SGD algorithm with one modification involving sampling multiple time steps of a single data point when computing the loss. Building on this, another study [18] further presented the effectiveness of three other techniques, namely pre-training, augmentation multiplicity, and modified time step sampling. While DP-SGD is deemed a strong countermeasure for privacy leaks, it comes at the cost of sample quality and longer training times.
3 Empirical Analysis